(square root of x+2)-x=0 Solve algebraically and check your potential solutions: I have no idea how to do this problem. Any suggestions will be very helpful
Answer by scott8148(6628)

adding x __ sqrt(x+2)=x

squaring __ x+2=x^2

subtracting x+2 __ 0=x^2-x-2

factoring __ 0=(x-2)(x+1)